# LightMem: Lightweight and Efficient Memory-Augmented Generation

### v1

@@ Appendix A Usage of LLMs
Appendix B Methodology Details
B.1 Topic Segmentation
@@
B.2 Category-wise Accuracy
@@
B.3 Detailed Parameter Analysis
@@
Appendix C Experiment Details
C.1 Datasets and Baselines
The LongMemEval dataset (Wu et al., 2025) is designed to benchmark long-term interactive memory in conversational agents. It comprises 500 evaluation questions built upon extended user-assistant dialogues. It has two different versions with different lengths: the LongMemEval-S setting contains approximately 115k tokens per problem, while the LongMemEval-M setting extends up to 1.5 million tokens across 500 sessions. In our work, we adopt the LongMemEval-S version due to its balance between dialogue length and computational feasibility. Each evaluation instance consists of a multi-session dialogue history, a question posed at a later timestamp, and the corresponding ground truth answer with supporting evidence spans. The questions are categorized into multiple types: information extraction, multi-session reasoning, knowledge updates, temporal reasoning, and abstention. Overall, the dataset is characterized by extremely long histories, wide temporal spans, and diverse question types, making it a comprehensive benchmark for evaluating conversational agents’ memory capabilities. During the experiments, five samples from this dataset contained corrupted characters, which caused LightMem’s compression model to fail to run properly. Consequently, LightMem directly discarded these five samples when processing the dataset. However, their accuracy results were uniformly treated as false. The indices of these five samples in the dataset are 74, 183, 278, 351, and 380.
@@ C.2 Implementation Details
Appendix D Prompts
D.1 LLM-as-Judge

### v3

@@ Appendix B Usage of LLMs
Appendix C Methodology Details
C.1 Topic Segmentation
@@
C.2 Category-wise Accuracy
@@
C.3 Detailed Parameter Analysis
@@
Table 8: The impact of LightMem compression ratio $r$ and STM buffer threshold $th$ is reported here. Due to space limitations, we only present a subset of representative results of the online soft update results, with more results provided in the Figure 9.
Model thth $r$ ACC Input (k) Output (k) Total (k) Calls Time
GPT
256 0.5 64.29 20.80 10.01 30.81 25.67 302.69
256 0.6 67.68 24.58 10.53 35.11 30.47 329.61
256 0.7 65.68 27.66 9.97 37.63 34.26 403.59
512 0.6 63.74 16.23 9.45 25.68 15.63 266.98
512 0.7 68.64 18.88 9.37 28.25 18.43 283.76
512 0.8 66.67 21.55 8.59 30.14 21.11 268.97
1024 0.6 59.68 10.34 7.68 18.20 7.69 177.45
1024 0.7 64.68 12.93 6.90 19.83 8.25 209.12
1024 0.8 64.35 14.86 6.28 21.14 9.43 216.08
Qwen
512 0.4 58.57 11.03 17.00 28.03 10.11 421.74
512 0.6 66.57 16.22 19.50 35.72 15.40 471.09
512 0.8 67.37 21.35 19.36 40.71 20.98 461.02
768 0.4 61.95 9.01 16.14 25.15 6.54 357.13
768 0.6 73.20 13.19 19.21 32.40 9.97 417.13
768 0.8 64.95 16.94 19.06 36.00 13.09 420.14
1024 0.4 53.91 8.02 15.44 23.46 4.83 300.56
1024 0.6 65.67 11.50 18.21 29.71 7.18 396.35
1024 0.8 68.69 14.82 18.49 33.31 9.43 355.71
@@
Appendix D Experiment Details
D.1 Datasets and Baselines
Datasets
The LongMemEval dataset (Wu et al., 2025) is designed to benchmark long-term interactive memory in conversational agents. It comprises 500 evaluation questions built upon extended user-assistant dialogues. It has two different versions with different lengths: the LongMemEval-S setting contains approximately 115k tokens per problem, while the LongMemEval-M setting extends up to 1.5 million tokens across 500 sessions. In our work, we adopt the LongMemEval-S version due to its balance between dialogue length and computational feasibility. The questions are categorized into multiple types: information extraction, multi-session reasoning, knowledge updates, temporal reasoning, and abstention. Overall, the dataset is characterized by extremely long histories, wide temporal spans, and diverse question types, making it a comprehensive benchmark for evaluating conversational agents’ memory capabilities. During the experiments, five samples from this dataset contained corrupted characters, which caused LightMem’s compression model to fail to run properly. Consequently, LightMem directly discarded these five samples when processing the dataset. However, their accuracy results were uniformly treated as false. The indices of these five samples in the dataset are 74, 183, 278, 351, and 380.
The LoCoMo benchmark targets the evaluation of long-range conversational memory. It features extremely long dialogues, with each conversation spanning roughly 300 turns and around 9K tokens on average. The accompanying questions fall into four categories—Single-hop, Multi-hop, Temporal, and Open-domain—providing a comprehensive assessment of different dimensions of memory in LLMs.
Baselines
@@ D.2 Implementation Details
Appendix E Prompts
E.1 LLM-as-Judge